OpenStudy (anonymous):

The monomial 36x4 is a perfect square. What is the square root of 36x4? 6x2 6x4 18x2 18x4

OpenStudy (anonymous):

@Nnesha Is It A?

OpenStudy (jhannybean):

what is 36 \(\cdot\) 4 =?

OpenStudy (jhannybean):

\[\sqrt{36 \cdot 4} = \sqrt{36} \cdot \sqrt{4} =~?\]

OpenStudy (jhannybean):

Can solve this a number of ways, actually.

OpenStudy (anonymous):

Is It A @Jhannybean

OpenStudy (jhannybean):

I won't tell you the answer. I want oyu to tell me what the square root of 36 is, and what the square root of 4 is.

OpenStudy (anonymous):

6 and 2

OpenStudy (jhannybean):

And what do you get when you multiply them?

OpenStudy (anonymous):

12

OpenStudy (jhannybean):

an what do you get when you square 12 ?

OpenStudy (michele_laino):

please note that \[36=6^{2}, 4=2^{2}\]

OpenStudy (anonymous):

4

OpenStudy (jhannybean):

no, what is \(12^2\)?

OpenStudy (anonymous):

144

OpenStudy (jhannybean):

Good. when you multiply 36 x 4 you get 144.

OpenStudy (jhannybean):

Therefore,from what you have told me so far, let's put it all together.

OpenStudy (anonymous):

6-4-12-144

OpenStudy (jhannybean):

\[36 \cdot 4 = 144 \implies \sqrt{144} = \sqrt{12 \cdot 12} = 12= 6 \cdot 2\]

OpenStudy (paxpolaris):

the square root of : \(36x^4\) is \(6x^2\) ... correct

OpenStudy (jhannybean):

Does that make sense, @iam.dhue ?

OpenStudy (anonymous):

Yea iHave It Written Down

OpenStudy (jhannybean):

Great :) so which answer did you get?

OpenStudy (anonymous):

A

OpenStudy (jhannybean):

Good job.

OpenStudy (anonymous):

You Mind Helping Me

OpenStudy (jhannybean):

Sure, I can try. close this question and open a new one :)